An electromagnetic wave at 100 Hz can penetrate to a depth of 100 m. Then, an electromagnetic wave of 400 Hz will penetrate to a depth of _________m.


1
Expert's answer
2021-02-25T11:23:12-0500

The penetration depth

"\delta=\sqrt{\frac{1}{\pi\mu f\sigma}}"

Hence,


"\delta_2=\delta_1\sqrt{\frac{f_1}{f_2}}\\
=100\:\rm m\sqrt{\frac{100\: Hz}{400\: Hz}}=50\:\rm m"